From 4 cards bearing the numbers -2, -1, 0, 1, one card will be drawn at random. Let X be the square of the number on the card. Find the expectaion of X. I have worked on this for hours but never seem to come up with a an answer that matches any of these: a 1, c 2 c 1.75 or d 1.5.
Answer by stanbon(48568)

From 4 cards bearing the numbers -2, -1, 0, 1, one card will be drawn at random. Let X be the square of the number on the card. Find the expectaion of X. I have worked on this for hours but never seem to come up with a an answer that matches any of these: a 1, c 2 c 1.75 or d 1.5.
-----------
Values of "X": 4, 1, 0, 1
Probability for each is (1/4)
----------------
Expected value of X: (1/4)(4+1+0+1) = 6/4 = 3/2 = 1.5
===========================================================
Cheers,
Stan H.